\begin{thebibliography}{99}  

\bibitem{CL22}
A. W. Cunha and E. L. Lima,
{\it A note on gradient k-Yamabe solitons},
Ann. Global Anal. Geom. {\bf 62} (2022), no. 1, 73-82.

\bibitem{BB13}
E. Barbosa and E. Ribeiro,
  {\it On conformal solutions of the Yamabe flow}, Arch. Math. {\bf 101} (2013), 79-89.

\bibitem{BBGG11}
W. Batat, M. Brozos-V{\'a}zquez, E. Garc{\'i}a-R{\'i}o and S. Gavino-Fern{\'a}ndez,
{\it Ricci solitons on Lorentzian manifolds with large isometry groups},
Bull. Lond. Math. Soc. {\bf 43} (2011), no. 6, 1219-1227.

\bibitem{BGG13a}
M. Brozos-V{\'a}zquez, E. Garc{\'i}a-R{\'i}o and S. Gavino-Fern{\'a}ndez,
{\it Locally conformally flat Lorentzian gradient Ricci solitons},
J. Geom. Anal. {\bf 23} (2013), no. 3, 1196-1212.

\bibitem{BGV16}
M. Brozos-V\'azquez, E. Garc\'ia-R\'io, and X. Valle-Regueiro,  
{\it Half conformally flat gradient Ricci almost solitons},
Proc. R. Soc. A \textbf{472}, 20160043, 12 pp.


\bibitem{Brendle}
S. Brendle, 
{\it Rotational symmetry of self-similar solutions to the Ricci flow},
Invent. Math. (2013) {\bf 194}, 731--764.

\bibitem{CC12}
H.-D. Cao and Q. Chen,
{\it On locally conformally flat gradient steady Ricci solitons},
Trans. Amer. Math. Soc. {\bf 364} (2012), 2377-2391.

\bibitem{CC13}
H.-D. Cao and Q. Chen, 
{\em On Bach-flat gradient shrinking Ricci solitons},
Duke Math. J.  (2013) {\bf 162}, 1149--1169.

\bibitem{CSZ12}
H.-D. Cao, X. Sun and Y. Zhang, 
{\it On the structure of gradient Yamabe solitons,}
Math. Res. Lett. (2012) {\bf 19}, 767--774.

\bibitem{CMM12}
G. Catino, C. Mantegazza and L. Mazzieri, 
{\it On the global structure of conformal gradient solitons with nonnegative Ricci tensor},
Commun. Contemp. Math. (2012) {\bf 14}, 12 pp.

\bibitem{CC96}
J. Cheeger and T. H. Colding, 
{\it Lower bounds on Ricci curvature and the almost rigidity of warped products,}
Ann. of Math. (1996) {\bf 144}, 189--237.


\bibitem{CMMR17}
G. Catino, P. Mastrolia, D. D. Monticelli and M. Rigoli,
{\it On the geometry of gradient Einstein-type manifolds},
Pacific J. Math. {\bf 286} (2017), no. 1, 39--67.

\bibitem{DS13}
P. Daskalopoulos and N. Sesum, 
\textit{The classification of locally conformally flat Yamabe solitons,}
Adv. Math. (2013) {\bf 240}, 346--369.

\bibitem{FG14}
M. Fern{\'a}ndez-L{\'o}pez and E. Garc{\'i}a-R{\'i}o,
{\it A note on locally conformally flat gradient Ricci solitons},
Geom. Dedicata {\bf 168} (2014), 1-7.

\bibitem{FM21}
S. Fujii and S. Maeta,
{\it Classification of generalized Yamabe solitons in Euclidean spaces},
Internat. J. Math. {\bf 32} (2021), 12 pages.

\bibitem{Hamilton89}
R. Hamilton, 
{\it Lectures on geometric flows},
(1989), unpublished.


\bibitem{HL14}
G. Huang and H. Li,
{\it On a classification of the quasi Yamabe gradient solitons},
Methods Appl. Anal. {\bf 21} (2014), no. 3, 379-390.

\bibitem{Hsu12}
S. Y. Hsu, 
\textit{A note on compact gradient Yamabe solitons,}
J. Math. Anal. Appl. (2012) {\bf 388} (2), 725--726.


\bibitem{ONiell}
B. O'Neill,
{\it Semi-Riemannian Geometry With Applications to Relativity,}
Academic Press, (1983).

\bibitem{Perelman1}
G. Perelman, 
{\it The entropy formula for the Ricci flow and its geometric applications,}
arXiv math.DG/0211159, (2002).

\bibitem{Tashiro65}
Y. Tashiro, {\it Complete Riemannian manifolds and some vector fields},
 Trans. Amer. Math. Soc., {\bf 117} (1965), 251--275.

\bibitem{Yano40}
K. Yano, {\it Concircular geometry I}, Proc. Imp. Acad. Tokyo, {\bf 16} (1940), 195--200.
 
\bibitem{Zeng20}
F. Zeng,
{\it On the h-almost Yamabe soliton},
J. Math. Study, {\bf 54} (2021), 371--386.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{latexsym, amssymb, amsmath}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{remark}[theorem]{Remark}
\theoremstyle{definition}

\makeatletter
    
    \@addtoreset{equation}{section}
  \makeatother

\newcommand{\Ric}{{\rm Ric}}

\newcommand{\n}{\nabla}

\begin{document}

\title[Classification of generalized Yamabe solitons]
{Classification of generalized Yamabe solitons under vanishing conditions on the Weyl, Cotton, and Cao-Chen tensors}

\author{Shun Maeta}
\address{Department of Mathematics,
Faculty of Education,
Chiba University, 1-33, Yayoicho, Inage, Chiba, Chiba, 263-8522, Japan.}
\curraddr{}
\email{shun.maeta@gmail.com~{\em or}~shun.maeta@chiba-u.jp}
\thanks{The author is partially supported by the 	
Grant-in-Aid for Scientific Research (C), No.23K03107, Japan Society for the Promotion of Science.\\
Department of Mathematics,
Chiba University, 1-33, Yayoicho, Inage, Chiba, 263-8522, Japan.\\
{\it Email address}: shun.maeta@gmail.com~{\em or}~shun.maeta@chiba-u.jp}

\subjclass[2010]{53C21, 53C25, 53C20}

\date{}

\dedicatory{}

\keywords{Gradient conformal solitons; Gradient Einstein-type manifolds; Yamabe solitons; locally conformally flat}

\commby{}

\begin{abstract}
We study complete conformal gradient solitons, a class containing gradient Yamabe solitons and many generalized Yamabe-type structures, including gradient almost Yamabe, gradient k-Yamabe, and gradient h-almost Yamabe solitons, and, after a change of the potential function, gradient Einstein-type manifolds with $\alpha=0$ and $\beta\neq0$ (in particular, quasi-Yamabe solitons). In this paper, we classify complete nontrivial locally conformally flat conformal gradient solitons. This result contributes to an analogue of Perelman's conjecture for Yamabe-type solitons. Moreover, we show that under nonnegative scalar curvature, every nonflat soliton is rotationally symmetric. We also obtain classifications assuming the Cotton or Cao-Chen tensor vanishes.

\end{abstract}

\maketitle

\section{Introduction}\label{intro} 

Let $(M,g)$ be an $n$-dimensional Riemannian manifold. For smooth functions $F$ and $\varphi$ on $M$, $(M,g,F,\varphi)$ is called a {\it conformal gradient soliton} (cf. \cite{Yano40}, \cite{Tashiro65}, \cite{CMM12}) if
\begin{equation}\label{GCS}
\varphi g=\nabla\nabla F.
\end{equation}
If $F$ is constant, then $(M,g,F,\varphi)$ is called trivial.
This class goes back to the work of Yano and Tashiro on concircular fields, and it also appears in the warped product rigidity theory of Cheeger and Colding \cite{CC96}.
In particular, \eqref{GCS} is a natural framework for studying global geometry through a conformal potential function.

A fundamental special case is the gradient Yamabe soliton
\begin{equation}\label{YS}
(R-\rho)g=\nabla\nabla F,
\end{equation}
where $R$ is the scalar curvature and $\rho\in\mathbb{R}$.
According to the sign of $\rho$, the soliton is called steady, shrinking, or expanding.
Gradient Yamabe solitons are self-similar solutions to the Yamabe flow introduced by Hamilton \cite{Hamilton89}, and they are expected to control singularity formation and asymptotic profiles of the flow.
It is also known that compact gradient Yamabe solitons are trivial, namely, $F$ is constant (see, for example, \cite{Hsu12}).

Equation \eqref{YS} is the scalar-curvature analogue of the gradient Ricci soliton equation.
For gradient Ricci solitons, Brendle \cite{Brendle} proved Perelman's conjecture \cite{Perelman1}, that is, rotational symmetry of three-dimensional complete noncompact $\kappa$-noncollapsed gradient steady Ricci solitons with positive curvature.
This motivates the Yamabe-soliton version of Perelman's conjecture, namely, to determine natural geometric assumptions under which complete nontrivial gradient Yamabe solitons must be rotationally symmetric.

In this direction, Daskalopoulos and Sesum \cite{DS13} proved rotational symmetry for complete locally conformally flat gradient Yamabe solitons with positive sectional curvature.
They also classified radial solutions and studied their asymptotic behavior.
Soon after that, Cao, Sun and Zhang \cite{CSZ12} established the global warped product structure of complete nontrivial gradient Yamabe solitons and classified complete locally conformally flat gradient Yamabe solitons.
In particular, in the locally conformally flat setting with $R\geq0$, their results imply rotational symmetry for nontrivial nonflat complete gradient Yamabe solitons.
From a broader viewpoint, Catino, Mantegazza and Mazzieri \cite{CMM12} studied conformal gradient solitons and proved a global structure theorem under nonnegative Ricci curvature.

The purpose of this paper is to unify and extend these Yamabe-type classification results within the conformal gradient soliton framework.
A key point is that \eqref{GCS} includes many generalized Yamabe-type structures by a suitable choice of $\varphi$.
This includes gradient almost Yamabe solitons, gradient $k$-Yamabe solitons, and gradient $h$-almost Yamabe solitons (see \cite{BB13}, \cite{CL22}, \cite{Zeng20}).
It also includes, in an essential sense, gradient Einstein-type manifolds with $\alpha=0$ and $\beta\not=0$.

More precisely, following Catino, Mastrolia, Monticelli and Rigoli \cite{CMMR17}, a gradient Einstein-type manifold is defined by
\[
\alpha \mathrm{Ric}+\beta \nabla\nabla f+\mu\, df\otimes df=(\varrho R+\lambda)g,
\]
where $\alpha,\beta,\mu,\varrho\in\mathbb{R}$ and $\lambda\in C^\infty(M)$.
When $\alpha=0$ and $\beta\not=0$, we have
\[
\beta \nabla\nabla f+\mu\, df\otimes df=(\varrho R+\lambda)g.
\]
If $\mu=0$, then $f$ itself is a conformal potential function up to a constant multiple.
If $\mu\not=0$ and we set $F=\frac{\beta}{\mu}e^{(\mu/\beta)f}$, then we have
\[
\nabla\nabla F=\frac{\mu}{\beta^2}(\varrho R+\lambda)F\, g.
\]
Hence, if we take $\varphi=\frac{\mu}{\beta^2}(\varrho R+\lambda)F$, the above equation is equivalent to \eqref{GCS}. Note that if $F$ is constant, then $f$ is also constant.
Therefore gradient Einstein-type manifolds with $\alpha=0$ and $\beta\not=0$ are essentially reduced to \eqref{GCS}.
In particular, this class contains quasi-Yamabe gradient solitons as a special case.

We remark that when $\alpha\not=0$, there are many studies assuming local conformal flatness and its generalizations (cf. \cite{BBGG11}, \cite{BGG13a}, \cite{BGV16}, \cite{CC13}, \cite{FG14}). 

Our main theorem gives a uniform rotational symmetry result for this broad class, including the above gradient Einstein-type manifolds, in the locally conformally flat case.

\begin{theorem}\label{thmA}
Any complete nontrivial nonflat locally conformally flat conformal gradient soliton with $R\geq0$ is rotationally symmetric.
\end{theorem}

Theorem~\ref{thmA} can be regarded as a generalized Yamabe-soliton analogue of the rotational symmetry results mentioned above.
As an immediate consequence, all the generalized Yamabe-type solitons listed above, as well as gradient Einstein-type manifolds with $\alpha=0$ and $\beta\not=0$ (including quasi-Yamabe gradient solitons), are rotationally symmetric under the assumptions of local conformal flatness and $R\geq0$.

We also study weaker conditions.
More precisely, we prove classification theorems under the vanishing of the Cotton tensor and the Cao-Chen tensor $D$ (introduced in \cite{CC13}, see also \cite{CC12}).
In the latter case, we obtain the following low-dimensional consequence.

\begin{theorem}\label{thmB}
For $n=3,4$, any complete nontrivial nonflat conformal gradient soliton $(M^n,g,F,\varphi)$ with $D\equiv0$ and $R\geq0$ is rotationally symmetric.
\end{theorem}

The paper is organized as follows.
In Section~\ref{Proof of main}, we provide a simpler proof of Tashiro's theorem. 
In Section~\ref{LCFGCS}, we classify complete conformal gradient solitons with $C\equiv0$. Using this classification, we classify locally conformally flat conformal gradient solitons and prove Theorem~\ref{thmA}.
In Section~\ref{vanish}, we study $D\equiv0$, and prove Theorem~\ref{thmB}.

\quad\\


\section{Lemmas for the main theorem}\label{Proof of main}
In this section, we provide a proof of a structure lemma for conformal gradient solitons.
We first recall some notation and definitions.
The Riemannian curvature tensor is defined by 
$$R(X,Y)Z=-\n_X\n_YZ+\n_Y\n_XZ+\n_{[X,Y]}Z,$$
for $X,Y,Z\in \mathfrak{X}(M)$. The Ricci tensor $R_{ij}$ is defined by 
$R_{ij}=R_{ipjp},$ where $R_{ijk\ell}=g(R({\partial_i,\partial_j})\partial_k,\partial_\ell).$

Conformal gradient solitons were studied by Cheeger and Colding \cite{CC96}. They gave a characterization of warped product manifolds. Inspired by their work, we will drastically simplify the proof of Tashiro's classification theorem \cite{Tashiro65}. In 2012, Catino, Mantegazza, and Mazzieri provided another proof of Tashiro's theorem \cite{CMM12}. We also provide a much simpler proof of Tashiro's theorem. 

\begin{lemma}\label{main}
A nontrivial complete conformal gradient soliton $(M^n,g,F,\varphi)$ is either

$(1)$ compact and rotationally symmetric, or

$(2)$ rotationally symmetric and equal to the warped product
$$([0,\infty),dr^2)\times_{|\n F|}(\mathbb{S}^{n-1},{\bar g}_{S}),$$
where $\bar g_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or

$(3)$ the warped product 
$$(\mathbb{R},dr^2)\times_{|\nabla F|} \left(N^{n-1},\bar g\right),$$
where the scalar curvature $\bar R$ of $N$ satisfies 
$$|\nabla F|^2R=\bar R-(n-1)(n-2)\varphi^2-2(n-1)g(\nabla F,\nabla\varphi).$$ 
\end{lemma}

\begin{remark}
By Lemma $\ref{main}$, to consider rotational symmetry of conformal gradient solitons, 
we only have to consider $(3)$ of Lemma $\ref{main}$.
\end{remark}

\begin{proof}[Proof of Lemma \ref{main}]

Let $c_0$ be a regular value of $F$, and $\Sigma_{c_0}=F^{-1}(c_0)$. Assume that $I(\ni c_0)$ is an open interval such that $F$ has no critical point in an open neighborhood $U_I=F^{-1}(I)$ of $\Sigma_{c_0}$. Then we have
$$g=\frac{1}{|\nabla F|^2}dF^2+\overline g_{\Sigma_{c_0}}=\frac{1}{|\nabla F|^2}dF^2+g_{ab}(F,x)dx^adx^b,$$
where $\overline g_{\Sigma_{c_0}}$ is the induced metric, $x=(x^2,\cdots,x^n)$ is a local coordinate system on $\Sigma_{c_0}$, and $a,b=2,3,\cdots,n.$ 

Since 
\begin{align*}
\nabla(|\nabla F|^2)=2\nabla\nabla F\nabla F=2\varphi g(\nabla F,\cdot),
\end{align*}
$|\nabla F|^2$ is constant on $\Sigma_{c}$ which is diffeomorphic to $\Sigma_{c_0}$.

On $U_I$, let $r=\int\frac{dF}{|\nabla F|}$. Then we have
$$g=dr^2+g_{ab}(r,x)dx^a d x^b.$$
Let $\nabla r:=\partial_1:=\partial_r\left(=\frac{\partial}{\partial r}\right)$. Then we have $|\nabla r|=1$ and $\nabla F=F'(r)\partial_1$. 
Without loss of generality, we may assume that $F'>0$ on $U_I$.
Assume that $J=(\alpha,\beta)$ with $F'(r)>0$ for all $r\in J$.
Since $\nabla_{\partial_1}\partial_1=0$, integral curves of $\nabla r$ are normal geodesics.
By the soliton equation, 
$$F''(r)=\varphi.$$
Thus, $\varphi$ is constant on $\Sigma_c.$
The second fundamental form can be written as 
$$B_{ab}=\frac{F''(r)}{F'(r)}g_{ab}.$$
Hence, the mean curvature can be written as $H=(n-1)\frac{F''(r)}{F'(r)}$.
By a direct computation, 
\begin{align*}
B_{ab}
=&g(\partial_1,-\nabla_a\partial_b)\\
=&-\Gamma_{ab}^1\\
=&-\frac{1}{2}g^{i\ell}\{\partial_ag_{\ell b}+\partial_bg_{a\ell}-\partial_\ell g_{ab}\}\\
=&\frac{1}{2}\partial_1g_{ab}.
\end{align*}
Thus, we have
$$\partial_1g_{ab}=2B_{ab}=2\frac{F''(r)}{F'(r)}g_{ab}.$$
Hence, we have
$$g_{ab}(r,x)=\left(\frac{F'(r)}{F'(r_0)}\right)^2g_{ab}(r_0,x).$$

Therefore, we have

$(1)$ $|\nabla F|$ and $\varphi$ are constant on $\Sigma_c$,

$(2)$ the second fundamental form of $\Sigma_c$ is $B_{ab}=\frac{\varphi}{|\nabla F|}g_{ab}$, 

$(3)$ the mean curvature $H=(n-1)\frac{\varphi}{|\nabla F|}$ is constant on $\Sigma_c$,

$(4)$ in any open neighborhood $F^{-1}((\alpha,\beta))$ of $\Sigma_c$ in which $F$ has no critical points, the soliton metric $g$ can be expressed as
$$g=dr^2+\frac{(F'(r))^2}{(F'(r_0))^2}\bar g_{r_0},$$
where $\bar g_{r_0}=g_{ab}(r_0,x)dx^adx^b$ is the induced metric on $\Sigma_c$, and $(x^2,\cdots,x^n)$ is a local coordinate system on $\Sigma_c$.  

The above argument shows that $|\nabla F|$ is constant on a regular level surface.
Set $N^{n-1}=F^{-1}(c_0)$ and $\overline g=\left(F'(r_0)\right)^{-2}\overline g_{r_0}$ for a regular value $c_0$ of $F$.
The local warped product expression obtained above extends along the integral curves of $\nabla r$ to a maximal interval $J$ on which $F'(r)=|\nabla F|>0$.
Since the warping function is $F'(r)$, every critical point of $F$ corresponds to a zero of $F'(r)$, and hence it can occur only at an endpoint of $J$.
Therefore, $F$ has at most two critical values.
After translating $r$, we may assume that one of the following holds:
$J=(-\infty,\infty),$
or
$
J=[0,\infty) \quad \text{with} \quad F'(0)=0,
$
or
$
J=[\alpha_0,\beta_0] \quad \text{with} \quad F'(\alpha_0)=F'(\beta_0)=0.
$

We consider the first case.  
By a direct calculation, we obtain the curvature formulas for the warped product with warping function $|\n F|=F'(r)>0$ (cf.~\cite{ONiell}). 

For $a,b,c,d=2,3,\cdots,n,$
\begin{align}\label{RT1}
R_{1a1b}&=-F'F'''{\bar g}_{ab},\quad R_{1abc}=0,\\
R_{abcd}&=(F')^2{\bar R}_{abcd}+(F'F'')^2(\bar g_{ad}\bar g_{bc}-\bar g_{ac}\bar g_{bd}),\notag
\end{align}
\begin{align}\label{RT2}
R_{11}=&-(n-1)\frac{F'''}{F'},\quad 
R_{1a}=0,\\
R_{ab}=&\bar R_{ab}-((n-2)(F'')^2+F'F''')\bar g_{ab},\notag
\end{align}
\begin{align}\label{RT3}
R=(F')^{-2}\bar R-(n-1)(n-2)\Big(\frac{F''}{F'}\Big)^2-2(n-1)\frac{F'''}{F'},
\end{align}
where the curvature tensors with bar are the curvature tensors of $(N,\bar g)$.

We consider the second case. Since $F$ has a unique critical point $x_0$, we have $r(x)={\rm dist}(x,x_0)$. Therefore, $\Sigma_c=\{F(x)=c\}$ is diffeomorphic to a geodesic sphere centered at $x_0$. By the smoothness of the metric $g$ at $x_0$, the induced metric $\overline g$ on $N^{n-1}$ is round. 

We consider the third case. It is immediate that $(M,g,F)$ is compact and rotationally symmetric.
\end{proof}

By Lemma $\ref{main}$, we recover the classification of three-dimensional complete gradient Yamabe solitons obtained in \cite{CSZ12}.
In fact, by the non-existence theorem for compact gradient Yamabe solitons (cf. \cite{Hsu12}), (1) of Lemma \ref{main} cannot happen. We will consider the case (3) of Lemma \ref{main}. 
By the soliton equation  $R-\rho=\varphi=F''$ and 
$$R=(F')^{-2}\bar R-2\Big(\frac{F''}{F'}\Big)^2-4\frac{F'''}{F'},$$
it follows that the scalar curvature $\bar R$ is constant. Therefore, $N^2$ is a space form.

\begin{corollary}\label{class3dim}
A nontrivial three-dimensional complete gradient Yamabe soliton $(M^3,g,F,\rho)$ is either

$(1)$ rotationally symmetric and equal to the warped product
$$([0,\infty),dr^2)\times_{|\n F|}(\mathbb{S}^{2},{\bar g}_{S}),$$
where $\bar g_{S}$ is the round metric on $\mathbb{S}^{2},$ or

$(2)$ the warped product 
$$(\mathbb{R},dr^2)\times_{|\nabla F|} \left(N^{2}(c),\bar g\right),$$
where $N^2(c)$ is a space form with constant curvature $c.$
\end{corollary}

Therefore, we give an affirmative partial answer to the Yamabe soliton version of Perelman's conjecture.

\begin{corollary}\label{PC0}
Any $3$-dimensional complete nontrivial nonflat gradient Yamabe soliton $(M^3,g,F,\rho)$ with $R\geq0$ is rotationally symmetric. 

\end{corollary}
\quad\\


\section{Proof of Theorem \ref{thmA}}\label{LCFGCS}

In this section, we prove Theorem \ref{thmA}. 
We first recall the Cotton tensor $C$ and the Weyl tensor $W$. 
\begin{align*}
C_{ijk}
=&\nabla_iS_{jk}-\nabla _jS_{ik}\\
=&\nabla_iR_{jk}-\nabla_jR_{ik}-\frac{1}{2(n-1)}(g_{jk}\nabla_iR-g_{ik}\nabla_jR),
\end{align*}
where $S=\Ric-\frac{1}{2(n-1)}Rg$ is the Schouten tensor.
 The Cotton tensor is skew-symmetric in the first two indices and totally trace-free, that is,
$$C_{ijk}=-C_{jik} \quad \text{and} \quad g^{jk}C_{ijk}=g^{ik}C_{ijk}=0.$$
\begin{align}\label{Weyl}
W_{ijk\ell}
=&R_{ijk\ell}-\frac{1}{n-2}(R_{ik}g_{j\ell}+R_{j\ell}g_{ik}-R_{i\ell}g_{jk}-R_{jk}g_{i\ell})\\
&+\frac{R}{(n-1)(n-2)}(g_{ik}g_{j\ell}-g_{i\ell}g_{jk}).\notag
\end{align}
As is well known, a Riemannian manifold $(M^n,g)$ is locally conformally flat if and only if 
(1) for $n\geq4$, the Weyl tensor vanishes; (2) for $n=3$, the Cotton tensor vanishes.
Moreover, for $n\geq4$, if the Weyl tensor vanishes, then the Cotton tensor vanishes. 
For $n=3$, the Weyl tensor always vanishes, whereas the Cotton tensor does not vanish in general.

First, we consider conformal gradient solitons under the weaker assumption that the Cotton tensor vanishes. This will be used to prove Theorem \ref{thmA}.

\begin{proposition}\label{subC}
A nontrivial complete conformal gradient soliton $(M^n,g,F,\varphi)$ with $C\equiv0$  is either

$(1)$ compact and rotationally symmetric, or

$(2)$ rotationally symmetric and equal to the warped product
$$([0,\infty),dr^2)\times_{|\n F|}(\mathbb{S}^{n-1},{\bar g}_{S}),$$
where $\bar g_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or

$(3)$ the warped product 
$$(\mathbb{R},dr^2)\times_{|\nabla F|} \left(N^{n-1},\bar g\right),$$
where $N$ has constant scalar curvature $\bar R$.
Furthermore, if $R\geq0$, then either $\bar R>0$, or $R=\bar R=0$ and $(M,g)$ is isometric to the Riemannian product $(\mathbb{R},dr^2)\times(N^{n-1},\bar g).$ 
\end{proposition}

\begin{proof}
We only have to consider the case (3) of Lemma \ref{main}.
By the same argument as in the proof of Lemma \ref{main}, we have
\begin{align}
R_{11}=&-(n-1)\frac{F'''}{F'},\quad 
R_{1a}=0,\\
R_{ab}=&\bar R_{ab}-((n-2)(F'')^2+F'F''')\bar g_{ab},\notag
\end{align}
\begin{align}\label{vvRT3}
R=(F')^{-2}\bar R-(n-1)(n-2)\Big(\frac{F''}{F'}\Big)^2-2(n-1)\frac{F'''}{F'},
\end{align}
for $a,b=2,3,\cdots,n.$
By the definition of the Cotton tensor, we have  
\begin{align*}
C_{1a1}
=&\nabla_1R_{a1}-\nabla_aR_{11}-\frac{1}{2(n-1)}(g_{a1}\nabla_1R-g_{11}\nabla_aR)\\
=&\frac{1}{2(n-1)}\nabla_a R.
\end{align*}
From this and \eqref{vvRT3}, the scalar curvature $\bar R$ of $N$ is constant.
Assume that $R\geq 0$. If $\bar R\leq0$, then by \eqref{vvRT3}, $F'''\leq0$. 
Therefore, $F'(>0)$ is concave, which means that $F'$ is constant. 
By \eqref{vvRT3} again, $R=\bar R=0$ and $(M,g)$ is the Riemannian product $(\mathbb{R},dr^2)\times(N^{n-1},\bar g).$
\end{proof}

We next show the following lemma.

\begin{lemma}\label{nocriticalpoint}
Let $(M,g,F,\varphi)$ be a nontrivial complete locally conformally flat conformal gradient soliton. Assume that $F$ has no critical point. Then, $(M,g,F,\varphi)$ is a warped product
$$(\mathbb{R},dr^2)\times_{|\nabla F|} \left(N^{n-1}(c),\bar g\right),$$
where $(N^{n-1}(c),\bar g)$ is a space form.
\end{lemma}

\begin{proof}
We consider $(3)$ of Lemma \ref{main}. By the same argument as in the proof of Lemma \ref{main}, one can obtain the curvature formulas for the warped product with warping function $|\n F|=F'(r)>0$. For $a,b,c,d=2,3,\cdots, n,$ 
\begin{align}\label{aRT1}
R_{1a1b}&=-F'F'''{\bar g}_{ab},\quad R_{1abc}=0,\\
R_{abcd}&=(F')^2{\bar R}_{abcd}+(F'F'')^2(\bar g_{ad}\bar g_{bc}-\bar g_{ac}\bar g_{bd}),\notag
\end{align}
\begin{align}\label{aRT2}
R_{11}=&-(n-1)\frac{F'''}{F'},\quad 
R_{1a}=0,\\
R_{ab}=&\bar R_{ab}-((n-2)(F'')^2+F'F''')\bar g_{ab},\notag
\end{align}
\begin{align}\label{aRT3}
R=(F')^{-2}\bar R-(n-1)(n-2)\Big(\frac{F''}{F'}\Big)^2-2(n-1)\frac{F'''}{F'}.
\end{align}

$\dim M=3$: By Proposition \ref{subC}, $N$ is a space form. 

$\dim M \geq4$:
By \eqref{Weyl}, \eqref{aRT1}, \eqref{aRT2} and \eqref{aRT3}, we have
\begin{align*}
W_{1a1b}=&-\frac{\bar R_{ab}}{n-2}+\frac{\bar R}{(n-1)(n-2)}\bar g_{ab},\\
W_{1abc}=&0,\\
W_{abcd}
=&(F')^2\Big(\bar W_{abcd}\\
&+\frac{1}{(n-2)(n-3)}
\Big\{\frac{2}{n-1}\bar R(\bar g_{ad}\bar g_{bc}-\bar g_{ac}\bar g_{bd})\\
&\hspace{100pt}
-(\bar R_{ad}\bar g_{bc}+\bar R_{bc}\bar g_{ad}-\bar R_{ac}\bar g_{bd}-\bar R_{bd}\bar g_{ac})\Big\}\Big).
\end{align*}
Since $M$ is locally conformally flat, we have
\begin{equation}\label{NisEin}
\bar R_{ab}=\frac{\bar R}{n-1}\bar g_{ab},
\end{equation}
and 
\begin{align}\label{NisLCF}
\bar W_{abcd}
=&-\frac{1}{(n-2)(n-3)}
\Big\{\frac{2}{n-1}\bar R(\bar g_{ad}\bar g_{bc}-\bar g_{ac}\bar g_{bd})\\
&\hspace{70pt}-(\bar R_{ad}\bar g_{bc}+\bar R_{bc}\bar g_{ad}-\bar R_{ac}\bar g_{bd}-\bar R_{bd}\bar g_{ac})\Big\}.\notag
\end{align}
Substituting \eqref{NisEin} into \eqref{NisLCF}, we have $\bar W_{abcd}=0.$
Therefore, $N$ is Einstein and locally conformally flat, which means that $N$ is a space form.
\end{proof}

Combining Lemma \ref{nocriticalpoint} with Lemma \ref{main}, we obtain the following.

\begin{theorem}\label{sub2}
A nontrivial complete locally conformally flat conformal gradient soliton $(M^n,g,F,\varphi)$ is either

$(1)$ compact and rotationally symmetric, or

$(2)$ rotationally symmetric and equal to the warped product
$$([0,\infty),dr^2)\times_{|\n F|}(\mathbb{S}^{n-1},{\bar g}_{S}),$$
where $\bar g_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or

$(3)$ the warped product 
$$(\mathbb{R},dr^2)\times_{|\nabla F|} \left(N^{n-1}(c),\bar g\right),$$
where $(N^{n-1}(c),\bar g)$ is a space form.
\end{theorem}

We are now ready to prove Theorem \ref{thmA}.
By Theorem \ref{sub2}, we only need to show $\bar R>0$ in (3) of Theorem \ref{sub2}. 
Assume that $R\geq0$. If $\bar R\leq0$, by \eqref{aRT3}, $F'''\leq0$ on $\mathbb{R}$.
Therefore, $F'(>0)$ is concave, which means that $F'$ is constant. 
By \eqref{aRT3} again, $R=\bar R=0$, which means that $M$ is flat. Therefore, $\bar R>0$, and we obtain Theorem \ref{thmA}.


\section{Gradient conformal solitons with a vanishing condition on the Cao-Chen tensor}\label{vanish}

For a gradient Ricci soliton satisfying $\Ric-\lambda g=\nabla\nabla F$, Cao and Chen introduced a new tensor $D$ (cf. \cite{CC13}, \cite{CC12}). We refer to this tensor as the {\it Cao-Chen tensor}. For $n\geq3$, 
\begin{align*}
D_{ijk}
=&\frac{1}{n-2}(R_{kj}\n_i F-R_{ki}\n_jF)\\
&+\frac{1}{(n-1)(n-2)} (R_{it}g_{jk}\n_t F-R_{jt}g_{ik}\n_t F)\\
&-\frac{R}{(n-1)(n-2)}(g_{kj}\n_i F-g_{ki}\n_j F).
\end{align*}
Roughly speaking, on gradient Ricci solitons, the Cao-Chen tensor estimates the difference between the Weyl tensor and the Cotton tensor, but an analogous interpretation does not hold for conformal gradient solitons.

Therefore, it is interesting to consider conformal gradient solitons with $D\equiv0$.

\begin{theorem}\label{main2}
A nontrivial complete conformal gradient soliton $(M^n,g,F,\varphi)$ with $D\equiv0$ is either

$(1)$ compact and rotationally symmetric, or

$(2)$ rotationally symmetric and equal to the warped product
$$([0,\infty),dr^2)\times_{|\n F|}(\mathbb{S}^{n-1},{\bar g}_{S}),$$
where $\bar g_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or

$(3)$ the warped product 
$$(\mathbb{R},dr^2)\times_{|\nabla F|} \left(N_{Ein}^{n-1},\bar g\right),$$
where $N_{Ein}$ is an Einstein manifold. 
Furthermore, if $R\geq0$, then either $\bar R>0$, or $R=\bar R=0$ and $(M,g)$ is isometric to the Riemannian product $(\mathbb{R},dr^2)\times(N^{n-1},\bar g)$, where $N$ is Ricci flat. 
\end{theorem}

\begin{proof}
We only have to consider (3) of Lemma \ref{main}. 
By the same argument as in the proof of Lemma \ref{main}, we have
\begin{align}\label{vRT2}
R_{11}=&-(n-1)\frac{F'''}{F'},\quad 
R_{1a}=0,\\
R_{ab}=&\bar R_{ab}-((n-2)(F'')^2+F'F''')\bar g_{ab},\notag
\end{align}
\begin{align}\label{vRT3}
R=(F')^{-2}\bar R-(n-1)(n-2)\Big(\frac{F''}{F'}\Big)^2-2(n-1)\frac{F'''}{F'},
\end{align}
for $a,b=2,3,\cdots,n.$
By the definition of the Cao-Chen tensor $D$, we have
\begin{align*}
D_{1ab}
=&\frac{1}{n-2}(R_{ba}\n_1 F-R_{b1}\n_aF)\\
&+\frac{1}{(n-1)(n-2)} (R_{1t}g_{ab}\n_t F-R_{at}g_{1b}\n_t F)\\
&-\frac{R}{(n-1)(n-2)}(g_{ba}\n_1 F-g_{b1}\n_a F)\\
=&\frac{1}{n-2}R_{ba}\n_1 F
+\frac{1}{(n-1)(n-2)} R_{11}g_{ab}\n_1 F\\
&-\frac{R}{(n-1)(n-2)}g_{ba}\n_1 F.
\end{align*}
Since the Cao-Chen tensor vanishes and $F'>0$, we have
$$R_{ab}=\frac{R-R_{11}}{n-1}g_{ab}.$$
From this, \eqref{vRT2} and \eqref{vRT3}, we have
$$\bar R_{ab}=\frac{\bar R}{n-1}\bar g_{ab}.$$
Hence, $N$ is an Einstein manifold. 
Thus, the scalar curvature $\bar R$ of $N$ is constant.
Assume that $R\geq 0$. If $\bar R\leq0$, then by \eqref{vRT3}, $F'''\leq0$. 
Therefore, $F'(>0)$ is concave, which means that $F'$ is constant. 
By \eqref{vRT3} again, $R=\bar R=0$ and $\overline{\rm Ric}=0$ on $N$. Therefore, $(M,g)$ is the Riemannian product $(\mathbb{R},dr^2)\times(N^{n-1},\bar g),$ where $N$ is Ricci flat.
\end{proof}

If $\dim N=n-1\leq3$, then the Einstein manifold $N$ is a space form. Hence, we obtain the following theorem.

\begin{theorem}
For $n=3,4$, a nontrivial complete conformal gradient soliton $(M^n,g,F,\varphi)$ with $D\equiv0$ is either

$(1)$ compact and rotationally symmetric, or

$(2)$ rotationally symmetric and equal to the warped product
$$([0,\infty),dr^2)\times_{|\n F|}(\mathbb{S}^{n-1},{\bar g}_{S}),$$
where $\bar g_{S}$ is the round metric on $\mathbb{S}^{n-1},$ or

$(3)$ the warped product 
$$(\mathbb{R},dr^2)\times_{|\nabla F|} \left(N^{n-1}(c),\bar g\right),$$
where $(N^{n-1}(c),\bar g)$ is a space form.
\end{theorem}

By the same argument as in Section \ref{LCFGCS}, we obtain Theorem \ref{thmB}.



\bibliographystyle{amsbook}

\end{document}